OpenStudy (mathlegend):

Write the Equation of a line. parallel to 2x-3y=6 and passing through (-2,5)

OpenStudy (mathlegend):

I know that parallel means the same slope.

OpenStudy (mathlegend):

However to get it parallel? I was thinking using y-y1=m(x-x1) ???

OpenStudy (anonymous):

yes, do that

hartnn (hartnn):

so the new line will also have same slope,just find m first and put (x1,y1)=(-2,5) in your equation,and u will get the equation of the line

OpenStudy (campbell_st):

write your given line in point slope form 3y = 2x - 6 y = 2/3 x - 2 now you have the slope of the line the general form is y = mx + b or in your case y = 2/3x + b substitute x and y from your point to find b

OpenStudy (mathlegend):

y=2x+9?

OpenStudy (campbell_st):

nope... not parallel

hartnn (hartnn):

what did u get b as?

OpenStudy (mathlegend):

Well this is what I did...

OpenStudy (anonymous):

slope aint 2

OpenStudy (mathlegend):

y-5=2(x+2)

OpenStudy (campbell_st):

rewrite your given line with y as the subject and in slope intercept form

OpenStudy (mathlegend):

oops I forgot to change it into slope intercept form

OpenStudy (campbell_st):

if you have 2x - 3y = 6... make 3y the subject...

OpenStudy (mathlegend):

Okay so I got ... y=2/3x-2

OpenStudy (campbell_st):

thats correct... what is the slope of the line

OpenStudy (mathlegend):

The slope is y= 2/3x + 19/3?

OpenStudy (campbell_st):

thats correct... well done

OpenStudy (mathlegend):

Thanks everyone!

OpenStudy (campbell_st):

the only thing you could do is write the equation in the general form 3y = 2x + 19 2x - 3y = -19